# A Closer Look at Time Steps is Worthy of Triple Speed-Up for Diffusion Model Training

### v2

@@ 3 Experiments
We first present visualizations in Sec. 3.1 and describe the main experiment setups in Sec. 3.2, including benchmark datasets, network architectures, training details, and evaluation protocols. In Sec. 3.3, we present the main results regarding both performance and efficiency. After that, in Sec. 3.6 and 3.4, we ablate the effectiveness of each designed module and evaluate the generalization of SpeeD in various diffusion tasks and settings. Finally, we investigate the compatibility of our approach with other recent methods in 3.5.
@@ 3.2 Implementation Details
Datasets. We mainly investigate the effectiveness of our approach on the following datasets: MetFaces (Karras et al., 2020) and FFHQ (Karras et al., 2019) are used to train unconditional image generation, CIFAR-10 (Krizhevsky et al., 2009) and ImageNet-1K (Deng et al., 2009) are used to train conditional image generation, and MS-COCO (Lin et al., 2014) is used to evaluate the generalization of our method in the text to image task. More details of these datasets can be found in the Appendix A.
Network architectures. U-Net (Ronneberger et al., 2015) and DiT (Peebles & Xie, 2023) are two famous architectures in the diffusion model area. We implement our approach on these two architectures and their variants. We follow the same hyper parameters as the baseline by default. More information about the details of the architectures can be found in Appendix A.3.
Training details. We train all models using AdamW  (Kingma & Ba, 2014; Loshchilov & Hutter, 2017) with a constant learning rate 1e-4. We set the maximum step in training to 1000 and use the linear variance. All images are augmented with horizontal flip transformations if not stated otherwise. Following common practice in the generative modeling literature, the exponential moving average (EMA) (Gardner Jr, 1985) of network weights is used with a decay of 0.9999. The results are reported using the EMA model. Details can be found in Tab. 7.
Evaluation protocols. In inference, we default to generating 10K images. Fréchet Inception Distance (FID) is used to evaluate both the fidelity and coverage of generated images.
@@
dataset MetFaces FFHQ iterations 10K 20K 30K 40K 50K 10K 20K 30K 40K 50K DiT-XL/2 (Peebles & Xie, 2023) 398.7 132.7 74.7 36.7 29.3 356.1 335.3 165.2 35.8 12.9 P2 (Choi et al., 2022) 377.1 328.2 111.0 27.3 23.3 368.7 357.9 158.6 35.5 15.0 Min-SNR (Hang et al., 2023) 389.4 313.9 52.1 31.3 28.6 376.3 334.1 151.2 34.0 12.6 Log-Normal (Esser et al., 2024) 311.8 165.1 63.9 51.1 47.3 307.6 165.1 63.9 51.1 47.3 CLTS (Xu et al., 2024) 375.0 57.2 28.6 24.6 23.5 336.1 329.1 173.4 33.7 12.7 SpeeD (ours) 367.3 23.4 22.6 22.1 21.1 322.1 320.0 91.8 19.8 9.9
@@
Before our comparison, we first introduce our baseline, i.e., DiT-XL/2, a strong image generation backbone as introduced in DiT (Peebles & Xie, 2023). We follow the hyperparameter settings from DiT and train DiT-XL/2 on MetFaces (Karras et al., 2020) and FFHQ (Karras et al., 2019), respectively. We compare our approach with two re-weighting methods: P2 (Choi et al., 2022) and Min-SNR (Hang et al., 2023), and two re-sampling methods: Log-Normal (Karras et al., 2022) and CLTS (Xu et al., 2024). In the evaluation, we use 10K generated images to calculate FID (Heusel et al., 2017) for comparison. To make a detailed comparison, all the approaches are trained for 50K iterations and we report the FID scores per 10K iterations.
As shown in Tab LABEL:tab:comparison, compared to DiT-XL/2, re-weighting, and re-sampling methods, our approach obtains the best FID results. Specifically, at the 50K iteration, compared to other methods, we reduce 2.3 and 2.6 FID scores on MetFaces and FFHQ at least. Another interesting finding is that the re-weighting methods reduce the FID very slowly at the beginning of the training, i.e., from the 10K to 20K iterations. That aligns with our analysis well: re-weighting methods involve a lot of steps from the convergence area. Based on the experimental results, the time steps that come from the convergence area indeed contribute limited to the diffusion training.
@@
In addition to the performance comparison, we also present the acceleration results of our SpeeD . This naturally raises a question: how to calculate the acceleration? Here, we follow the previous diffusion acceleration methods (Gao et al., 2023) and other efficient training papers (Qin et al., 2023; Xu et al., 2024): visualizing the FID-Iteration curve and reporting the estimated highest acceleration ratio. We mainly compare with DiT-XL/2, one re-weighting method Min-SNR (Choi et al., 2022), and one re-sampling method CLTS (Xu et al., 2024) in Fig 5. At the same training iterations, our approach achieves significantly better FID scores than other methods. Notably, SpeeD accelerates the Min-SNR, and CLTS by 2.7 and 2.6 times, respectively. More comparisons with other methods can be found in Appendix B.
@@
Figure 5: We plot the FID-Iteration curves of our approach and other methods on the MetFaces dataset. SpeeD accelerates other methods obviously. The horizontal and vertical axes represent the training iterations and log(FID), respectively.
@@ 3.4 Generalization Evaluation
Cross-architecture robustness evaluation. There are mainly two architectures in the diffusion models: U-Net (Ronneberger et al., 2015) and DiT (Peebles & Xie, 2023). SpeeD is not correlated to specific model architecture, thereby it is a model-agnostic approach. We implement our method with DiT-XL/2 and U-Net on MetFaces, FFHQ, and ImageNet-1K, respectively. We default to training the models for 50K iterations on MetFaces and FFHQ, 400K on ImageNet-1K.To ensure a fair comparison, we keep all hyper-parameters the same and report the FID scores at 50K iterations. As shown in Tab. 4, SpeeD consistently achieve significantly higher performance under all settings, which indicates the strong generality of our method for different architectures and datasets.
Cross-schedule robustness evaluation. In the diffusion process, there are various time step schedules, including linear (Ho et al., 2020), quadratic and cosine (Nichol & Dhariwal, 2021) schedules. We verify SpeeD’s robustness across these schedules. We report FID and inception score (IS) (Salimans et al., 2016) scores as metrics for comparisons. As shown in Tab. 4, SpeeD achieves significant improvement on linear, quadratic, and cosine schedules both in FID and IS. That shows the strong generality of SpeeD in various schedules.
@@
[…]
Cross-task robustness evaluation. We apply SpeeD to the text-to-image task for evaluating the generality of our method. For text-to-image generation, we first introduce CLIP (Radford et al., 2021) to extract the text embedding for MS-COCO (Lin et al., 2014) dataset. Then, DiT-XL/2 is used to train a text-to-image model as our baseline. Following prior work(Saharia et al., 2022), FID score and CLIP score on MS-COCO validation set are evaluation metrics for quantitative analyses. As illustrated in Tab. 5, we obtain the better FID and CLIP score than our baseline.
@@ 3.5 Compatibility with other acceleration methods
Until now, we evaluate the effectiveness and generalization of our proposed method: SpeeD is a task-agnostic and architecture-agnostic diffusion acceleration approach. Is SpeeD compatible with other acceleration techniques? To investigate this, we apply our approach with two recent proposed algorithms: masked diffusion transformer (MDT) (Gao et al., 2023) and fast diffusion model (FDM) (Wu et al., 2023).
MDT + SpeeD. MDT (Gao et al., 2023) proposes a masked diffusion transformer method, which applies a masking scheme in latent space to enhance the contextual learning ability of diffusion probabilistic models explicitly. MDT can speed up the diffusion training by 10 times. They evaluate their MDT with DiT-S/2. We just inject our SpeeD on their MDT and report FID-Iteration curves for comparison in Fig. 6(a). All the results are obtained on ImageNet-1K dataset. Our approach can further accelerate MDT at least by 4 $\times$, which indicates the good compatibility of SpeeD.
FDM + SpeeD. Fast Diffusion Model (Wu et al., 2023) is a diffusion process acceleration method inspired by the classic momentum approach to solve optimization problem in parameter space. By integrating momentum into the diffusion process, it achieves similar performance as EDM (Karras et al., 2022) with less training cost. Based on the official implementation, we compare EDM, FDM, and FDM + SpeeD on CIFAR-10 of $32\times 32$ images. FDM accelerates the EDM by about 1.6 $\times$. Meanwhile, SpeeD can further reduce the overall training cost around 1.6 $\times$.
@@ 3.6 Ablation Experiments
We perform extensive ablation studies to illustrate the characteristics of SpeeD . The experiments in ablation studies are conducted on the FFHQ dataset and U-Net model by default. We ablate our designed components: asymmetric sampling (abbreviated as asymmetric) and change-aware weighting (abbreviated as CAW), suppression intensity $k$ in asymmetric sampling defined in Eqn. 3 and symmetry ceiling $\lambda$ for weighting in Sec. 2.6.
Table 6: Ablation studies on FFHQ dataset. Default settings and baseline are in gray and cyan.
@@
The symmetry ceiling $\lambda$ is a hyper-parameter that regulates the curvature of the weighting function. $\lambda$ is set in the interval $[0.5,1]$. The midpoint of the re-scaled weight interval is fixed at $0.5$. The symmetry ceiling $\lambda$ is the right boundary of the interval and the left boundary is $1$-$\lambda$. A higher $\lambda$ results in a larger weight interval and a more obvious distinction in weights between different time steps. In Tab. 6(c), settings $\lambda\leq 0.8$ obtain higher performance on FID scores than the baseline, which indicates SpeeD is relatively robust on symmetry ceiling $\lambda$. Further increase $\lambda$ leads to performance degradation. This means weighting should be in moderation.
@@ 4 Related Work
We discuss the related works about Diffusion Models and its Training Acceleration. The most related works are as follows, and more about cross-modality and video generation are in Appendix C.
@@
Diffusion models have emerged as the dominant approach in generative tasks (Saharia et al., 2022; Chen et al., 2023b; Wang et al., 2024a; Igashov et al., 2024), which outperform other generative methods including GANs (Zhu et al., 2017; Isola et al., 2017; Brock et al., 2018), VAEs (Kingma & Welling, 2013), flow-based models (Dinh et al., 2014). These methods (Ho et al., 2020; Song et al., 2020; Karras et al., 2022) are based on non-equilibrium thermodynamics (Jarzynski, 1997; Sohl-Dickstein et al., 2015), where the generative process is modeled as a reverse diffusion process that gradually constructs the sample from a noise distribution (Sohl-Dickstein et al., 2015). Previous works focused on enhancing diffusion models’ generation quality and alignment with users in visual generation. To generate high-resolution images, Latent Diffusion Models (LDMs) (Saharia et al., 2022; Rombach et al., 2022) perform diffusion process in latent space instead of pixel space, which employ VAEs to be encoder and decoder for latent representations.
@@
To reduce the computational costs, previous works accelerate the diffusion models in training and inference. For training acceleration, the earliest works (Choi et al., 2022; Hang et al., 2023) assign different weights to each time step on Mean Square Error (MSE) loss to improve the learning efficiency. The other methods in training acceleration are proposed, e.g., network architecture (Ryu & Ye, 2022; Wang et al., 2024b) and diffusion algorithm (Karras et al., 2022; Wu et al., 2023). Masking modeling (Gao et al., 2023; Zheng et al., 2023) are recently proposed to train diffusion models. Works  (Go et al., 2023; Park et al., 2024b; a; Kim et al., 2024) provide observations for explanation from the perspective of multi-tasks learning. SpeeD is of good compatibility with these methods, e.g., (Yu et al., 2024; Zheng et al., 2024a; b). In the field of sampling acceleration, a number of works tackle the slow inference speed of diffusion models by using fewer reverse steps while maintaining sample quality, including DDIM (Song et al., 2020), Analytic-DPM (Bao et al., 2022), and DPM-Solver (Lu et al., 2022).
@@
To better control the generation process with various conditions, e.g., image style, text prompt and stroke, Classifier-Free Guidance (CFG) proposes a guidance strategy with diffusion models that balance the sample quality and prompt alignment. ControlNet (Zhang et al., 2023b) reuses the large-scale pre-trained layers of source models to build a deep and strong encoder to learn specific conditions. Recently benefiting from diffusion models in the image generation field, video generation (Ma et al., 2024; Lab & etc., 2024) is getting trendy.
@@ 5 Conclusion
We present SpeeD , a novel approach for accelerating diffusion training by closely examining time steps. The core insights from this examination are: 1) suppressing the sampling probabilities of time steps that offer limited benefits to diffusion training (i.e., those with extremely small losses), and 2) emphasizing the importance of time steps with rapidly changing process increments. SpeeD demonstrates strong robustness across various architectures and datasets, achieving significant acceleration on multiple diffusion-based image generation tasks. Additionally, SpeeD is easily compatible with other diffusion acceleration methods, highlighting its wide applicability. We provide extensive theoretical analysis in this paper, aiming to support future research in both academia and industry.
@@ Acknowledgments.
This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-PhD-2021-08-008). This work is partially supported by the National Natural Science Foundation of China (62176165), the Stable Support Projects for Shenzhen Higher Education Institutions (20220718110918001), the Natural Science Foundation of Top Talent of SZTU(GDRC202131). Yang You’s research group is being sponsored by NUS startup grant (Presidential Young Professorship), Singapore MOE Tier-1 grant, ByteDance grant, ARCTIC grant, SMI grant (WBS number: A-8001104-00-00), Alibaba grant, and Google grant for TPU usage. We thank Tianyi Li, Yuchen Zhang, Yuxin Li, Zhaoyang Zeng, and Yanqing Liu for the comments on this work. Kai Wang (idea, writing, story, presentation), Yukun Zhou (implementation), and Mingjia Shi (theory, writing, presentation) contribute equally to this work. Xiaojiang Peng, Hanwang Zhang, and Yang You are equal advising. Xiaojiang Peng is the corresponding author.

### v3

@@ 3 Experiments
In this section, the visualization is provided in Section 3.1. Comparison with SOTA methods in is provided mainly on ImageNet-1K dataset. The compared baselines and scheduler settings including SDE-VP/VE and EDM. The comparison and ablations on most related hyperparameters and tasks are empirically verified.
@@ 3.2 Implementation Details
Datasets. We mainly investigate the effectiveness of our approach on the following datasets: MetFaces [27] and FFHQ [26] are used for unconditional tasks, Celeb-A [38], CIFAR-10 [33] and ImageNet-1K [9] are used to train conditional image generation, and MS-COCO [35] is used to evaluate the generalization of our method in the text to image task. More details can be found in the Appendix A.
Network architectures. U-Net [52] and DiT [47] are two famous architectures in the diffusion model area. We implement our approach on these two architectures and their variants. We follow the same hyper parameters as the baseline by default. More information about the details of the architectures can be found in Appendix A.1.
Training details. We train all models using AdamW  [31, 39] with a constant learning rate 1e-4. We set the maximum step in training to 1000 and use the linear variance. All images are augmented with horizontal flip transformations if not stated otherwise. Following common practice in the generative modeling literature, the exponential moving average (EMA) [13] of network weights is used with a decay of 0.9999. The results are reported using the EMA model. Details can be found in Tab. 6.
Evaluation protocols. In inference, we default to generating 10K images. Fréchet Inception Distance (FID) is to evaluate both the fidelity and coverage of generated images.
@@
dataset MetFaces FFHQ
iterations 10K 20K 30K 40K 50K 10K 20K 30K 40K 50K
DiT-XL/2 [47] 398.7 132.7 74.7 36.7 29.3 356.1 335.3 165.2 35.8 12.9
P2 [8] 377.1 328.2 111.0 27.3 23.3 368.7 357.9 158.6 35.5 15.0
Min-SNR [17] 389.4 313.9 52.1 31.3 28.6 376.3 334.1 151.2 34.0 12.6
Log-Normal [11] 311.8 165.1 63.9 51.1 47.3 307.6 165.1 63.9 51.1 47.3
CLTS [67] 375.0 57.2 28.6 24.6 23.5 336.1 329.1 173.4 33.7 12.7
SpeeD (ours) 367.3 23.4 22.6 22.1 21.1 322.1 320.0 91.8 19.8 9.9
@@
Before our comparison, we first introduce our baseline, i.e., DiT-XL/2, a strong image generation backbone as introduced in DiT [47]. We follow the hyperparameter settings from DiT and train DiT-XL/2 on MetFaces [27] and FFHQ [26], respectively. We compare our approach with two re-weighting methods: P2 [8] and Min-SNR [17], and two re-sampling methods: Log-Normal [28] and CLTS [67]. In the evaluation, we use 10K generated images to calculate FID [19] for comparison. To make a detailed comparison, all approaches are trained for 50K iterations and we report the FID per 10K iterations.
As shown in Tab 1, compared to DiT-XL/2, re-weighting, and re-sampling methods, our approach obtains the best FID results. Specifically, at the 50K iteration, compared to other methods, we reduce 2.3 and 2.6 FID scores on MetFaces and FFHQ at least. Another interesting finding is that the re-weighting methods reduce the FID slowly at the beginning of the training, i.e., from the 10K to 20K iterations. That aligns with our analysis well: re-weighting methods involve a lot from the convergence area. Based on the experimental results, the time steps from the convergence area indeed contribute limited to training.
@@
In addition to the performance comparison, we also present the acceleration results of our SpeeD . This naturally raises a question: how to calculate the acceleration? Here, we follow the previous diffusion acceleration methods [12] and other efficient training papers [48, 67]: visualizing the FID-Iteration curve and reporting the estimated highest acceleration ratio. We mainly compare with DiT-XL/2, one re-weighting method Min-SNR [8], and one re-sampling method CLTS [67] in Fig 6. At the same training iterations, our approach achieves significantly better FID scores than other methods. Notably, SpeeD accelerates the Min-SNR, and CLTS by 2.7 and 2.6 times, respectively (More comparisons in Appendix B).
@@
Figure 6: We plot the FID-Iteration curves of our approach and other methods on the MetFaces dataset. SpeeD accelerates other methods obviously. The horizontal and vertical axes represent the training iterations and log(FID), respectively. Detailed ones are in appendix.
@@ 3.4 Generalization Evaluation
Cross-architecture robustness evaluation. There are mainly two architectures in the diffusion models: U-Net [52] and DiT [47]. SpeeD is not correlated to specific model architecture, thereby it is a model-agnostic approach. We implement our method with DiT-XL/2 and U-Net on MetFaces, FFHQ, and ImageNet-1K, respectively. We default to training the models for 50K iterations on MetFaces and FFHQ, 400K on ImageNet-1K.To ensure a fair comparison, we keep all hyper-parameters the same and report the FID scores at 50K iterations. As shown in Tab. 2, SpeeD consistently achieve significantly higher performance under all settings, which indicates the strong generality of SpeeD for different architectures and datasets.
Cross-schedule robustness evaluation. In the diffusion process, there are various time step schedules, including linear [20], quadratic and cosine [42] schedules. We verify SpeeD’s robustness across these schedules. We report FID and inception score (IS) [55] scores as metrics. As shown in Tab. 3, SpeeD achieves significant improvement on linear, quadratic, and cosine schedules both in FID and IS, showing the generality of SpeeD in various schedules.
Cross-task robustness evaluation. We apply SpeeD to the text-to-image task for evaluating the generality of our method. For text-to-image generation, we first introduce CLIP [49] to extract the text embedding for MS-COCO [35] dataset. Then, DiT-XL/2 is used to train a text-to-image model as our baseline. Following prior work[54], FID score and CLIP score on MS-COCO validation set are
[…]
baseline 27.41 0.237
SpeeD 25.30 0.244
@@
evaluation metrics for quantitative analyses. As illustrated in Tab. 4, we obtain the better FID and CLIP score than our baseline.
@@ 3.5 Compatibility with other acceleration methods
Until now, we evaluate the effectiveness and generalization of our proposed method: SpeeD is a task-agnostic and architecture-agnostic diffusion acceleration approach. Is SpeeD compatible with other acceleration techniques? To investigate this, we apply our approach with two recent proposed algorithms: masked diffusion transformer (MDT) [12] and fast diffusion model (FDM) [66].
MDT + SpeeD. MDT [12] proposes a masked diffusion transformer method, which applies a masking scheme in latent space to enhance the contextual learning ability of diffusion probabilistic models explicitly. MDT can speed up the diffusion training by 10 times. They evaluate their MDT with DiT-S/2. We just inject our SpeeD on their MDT and report FID-Iteration curves for comparison in Fig. 7(a). All the results are obtained on ImageNet-1K dataset. Our approach can further accelerate MDT at least by 4 $\times$, which indicates the good compatibility of SpeeD.
FDM + SpeeD. Fast Diffusion Model [66] is a diffusion process acceleration method inspired by the classic momentum approach to solve optimization problem in parameter space. By integrating momentum into the diffusion process, it achieves similar performance as EDM [28] with less training cost. Based on the official implementation, we compare EDM, FDM, and FDM + SpeeD on CIFAR-10 of $32\times 32$ images. FDM accelerates the EDM by about 1.6 $\times$. SpeeD can further reduce the training cost around 1.6 $\times$.
@@ 3.6 Ablation Experiments
We perform extensive ablation studies to illustrate the characteristics of SpeeD . The experiments in ablation studies are conducted on the FFHQ dataset and U-Net model by default. We ablate our designed components: asymmetric sampling (abbreviated as asymmetric) and change-aware weighting (abbreviated as CAW), suppression intensity $k$ in asymmetric sampling defined in Eqn. 2 and symmetry ceiling $\lambda$ for weighting in Sec. 2.4.
@@
Table 5: Ablation studies on FFHQ dataset. Default settings and baseline are in purple and gray.
@@
The symmetry ceiling $\lambda$ is a hyper-parameter that regulates the curvature of the weighting function. $\lambda$ is set in the interval $[0.5,1]$. The midpoint of the re-scaled weight interval is fixed at $0.5$. The symmetry ceiling $\lambda$ is the right boundary of the interval and the left boundary is $1$-$\lambda$. A higher $\lambda$ results in a larger weight interval and a more obvious distinction in weights between different time steps. In Tab. 5(c), settings $\lambda\leq 0.8$ obtain higher performance on FID scores than the baseline, which indicates SpeeD is relatively robust on symmetry ceiling $\lambda$. Further increase $\lambda$ leads to performance degradation, and Weighting should be in moderation.
@@ 4 Related Work
We discuss the related works about Diffusion Models and its Training Acceleration. The most related works are as follows, and more discussion are in Appendix C.
@@
Diffusion models have emerged as the dominant approach in generative tasks [54, 7, 62, 23], which outperform other generative methods including GANs [82, 24, 4], VAEs [32], flow-based models [10]. These methods [20, 58, 28] are based on non-equilibrium thermodynamics [25, 57], where the generative process is modeled as a reverse diffusion process that gradually constructs the sample from a noise distribution [57]. Previous works focused on enhancing diffusion models’ generation quality and alignment with users in visual generation. To generate high-resolution images, Latent Diffusion Models (LDMs) [54, 51] perform diffusion process in latent space instead of pixel space, which employ VAEs to be encoder and decoder for latent representations.
@@
To reduce the computational costs, previous works accelerate the diffusion models in training and inference. For training acceleration, the earliest works [8, 17] assign different weights to each time step on Mean Square Error (MSE) loss to improve the learning efficiency. The other methods in training acceleration are proposed, e.g., network architecture [53, 64] and diffusion algorithm [28, 66]. Masking modeling [12, 74] are recently proposed to train diffusion models. Works  [14, 46, 45, 30] provide observations for explanation from the perspective of multi-tasks learning. SpeeD is of good compatibility with these methods, e.g., [69, 78, 79]. In the field of sampling acceleration, a number of works tackle the slow inference speed of diffusion models by using fewer reverse steps while maintaining sample quality, including DDIM [58], Analytic-DPM [1], and DPM-Solver [40].
@@
To better control the generation process with various conditions, e.g., image style, text prompt and stroke, Classifier-Free Guidance (CFG) proposes a guidance strategy with diffusion models that balance the sample quality and prompt alignment. ControlNet [71] reuses the large-scale pre-trained layers of source models to build a deep and strong encoder to learn specific conditions. Recently benefiting from diffusion models in the image generation field, video generation is getting trendy. The promising results are provided in recent works [41, 34] on video generation tasks.
@@ 5 Conclusion
SpeeD , a approach for accelerating diffusion training by closely examining time steps is proposed. The core insights are: 1) suppressing the sampling probabilities of time steps that offer limited benefits to diffusion training (i.e., those with extremely small losses), and 2) emphasizing the importance of time steps with rapidly changing process increments. SpeeD demonstrates strong robustness across various architectures and datasets, achieving significant acceleration on multiple diffusion-based image generation tasks. Extensive theoretical analyses in this paper, are provided.
Referring to previous works [47, 14, 46, 8, 17, 11, 67], our experiments train 50K, 100K, and 400K in MetFaces 256 × 256, FFHQ 256 × 256, and ImageNet-1K, respectively. 7M iterations are used in DiT to train ImageNet in the original work [47], which 1) is not the amount of resource usage that can be achieved in general research. Meanwhile, 2) the focus is on the acceleration effect of training, and a direct comparison of the ultimate convergence stage is unnecessary for this work.
@@ Acknowledgments.
This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-PhD-2021-08-008). This work is partially supported by the National Natural Science Foundation of China (62176165), the Stable Support Projects for Shenzhen Higher Education Institutions (20220718110918001), the Natural Science Foundation of Top Talent of SZTU(GDRC202131). Yang You’s research group is being sponsored by NUS startup grant (Presidential Young Professorship), Singapore MOE Tier-1 grant, ByteDance grant, ARCTIC grant, SMI grant (WBS number: A-8001104-00-00), Alibaba grant, and Google grant for TPU usage. We thank Tianyi Li, Yuchen Zhang, Yuxin Li, Zhaoyang Zeng, and Yanqing Liu for the comments on this work. Xiaojiang Peng, Hanwang Zhang, and Yang You are equal advising. Xiaojiang Peng is the corresponding author.
Supplementary Material